I am trying to figure out how to work this radical equation. 3squarerootx-5=3.

Answer by stanbon(48535)

3sqrt(x-5) = 3
Divide both sides by 3 to get:

sqrt(x-5) = 1
Square both sides to get:
x-5 = 1
x = 6
----------
Checking:
If x = 6
3sqrt(x-5) = 3sqrt(6-5) = 3sqrt(1) = 3*1 = 3
Correct.
=================
Cheers,
Stan H.